A sample of 900 items is found to have a mean of 3.47 cm. Write down the R programming code to test whether it can be reasonably regarded as a simple sample from a population with mean 3.23 cm and SD 2.31 cm at 99% level of confidence.


1
Expert's answer
2021-05-07T11:33:04-0400

"H_0 : \mu = 3.23 \\

H_1 : \mu ≠ 3.23 \\

\bar{x} = 3.47 \\

\sigma = 2.31 \\

n = 900"

The test statistic:

"z = \frac{\bar{x} - \mu}{\sigma / \sqrt{n}} \\

z = \frac{3.47 -3.23}{2.31 / \sqrt{900}} \\

z = 3.11"

R programming code:



Calculated "z>z_{alpha}" , so H0 is rejected.